Q.in a gravity free space,a man of mass "M" standing at a height "h" above the ground throws a ball of mass "m" straight down with speed "u".when the ball reaches the ground the distance of the man from above the ground is?


ANS::


h(1+m/M)


HELP!!!

Grade:12

5 Answers

Chetan Mandayam Nayakar
13 years ago

let speed of man be v. from principle of conservation of momentum, mu = Mv, v=mu/M, let time taken to reach ground be 't'. ut=h, t=h/u. therefore man moves vt=vh/u=mh/M, therefore, answer=h+(mh/M) =h(1+(m/M))

'Win exciting gifts by answering the questions on Discussion Forum.

harsha 9128
13 years ago

thanks for the replies!!!

if the speed of man is"v" shouldnt the conservation of momentum be as follows

(M+m)v=mu....

or is it only applicable when the velocity of ball is "u" relative to man?

thanks again.

Chetan Mandayam Nayakar
13 years ago

both u and v are relative to ground

PIYUSH AGRAWAL
13 years ago

by conservation of momentum in vertical direction,

0=mu+Mv

or, v=-mu/M

so, velocity of man is mu/M in upward direction.

time taken by ball to reach ground,as acc. is zero, t=h/u

in this time, man moves up by a distance s=vt=mh/M

so, man is at height s+h=(1+m/M)h above the ground.

 

ask more....

piyush agrawal

SIDDHI AMRALE
5 years ago
Here there is no external unbalanced force, so the law of conservation of momentum can be applied.
Pi=Pf
Initially, man at rest.
Pi= 0
When he throws the ball, the ball will hit the ground but as there is no gravity, the man will fly upwards.
Pf=Mv+(m (-u))
Where M is the mass of man, v is the velocity of man
m is the mass of ball, u is the velocity of ball
0= Mv+(-mu)
Mv=my
V= mu÷M      ( 1)
Consider after the ball is thrown, the man is flown up by distance x
As velocity is distance by time
V= x/t             (2)
From 1 and 2
mu/M=x/t
x= (mu/M) ×t
For ball,
u= h/t
X= mu/t  ×  h/u
So man will be at a distance of h+ x
h+x= h+ mh/M
       = h (1+m/M)